Starting in 2013, every guest staying at a Disney hotel is given a free magic band (a bracelet that has a computer chip that turns a bracelet into a hotel room key, parktickets, and dining plan meal card). Prof. K has a collection of 21 magic band bracelets from various trips over the years as listed below:
MAGIC BAND COLOR
NUMBER OF MAGIC BANDS
Red
7
Pink
1
Yellow
3
Blue
4
Orange
2
Purple
3
Pink
1
Gray
1
I. If perso grabs 3 random magic bands from the box containing her collection to wear on her next trip, what is the probability that she packs 3 red bands?
shhow all work
Answer #1

here total number of magic bands =21

number of red magic bands =7

therefore probability that she packs 3 red bands =P(first is red band)*P(2nd is red|first is red)*P(3rd is red|first 2 are red)

=(7/21)*(6/20)*(5/19)=210/7980 =21/798=1/38